Please help. I am so lost.
A cellular phone company offers a contract for which the cost C, in dollars, of t minutes of telephoning is given by C=0.25(t-500)+ 56.95, where it is assumed that t is greater than or equal to 500 minutes. What times will keep cost between $94.95 and $144.45?
For the cost to be between $94.95 and $144.45, the telephoning time must be between ______ minutes and _________minutes.

Answer by nerdybill(7384)

A cellular phone company offers a contract for which the cost C, in dollars, of t minutes of telephoning is given by C=0.25(t-500)+ 56.95, where it is assumed that t is greater than or equal to 500 minutes. What times will keep cost between $94.95 and $144.45?
.
The problem provides the formula for cost:
C=0.25(t-500)+ 56.95
where
C is cost
t is minutes
.
They want to know the "minutes" used if the cost is between $94.95 and $144.45
So, plug each cost and solve for t.
.
Lower bound:
C=0.25(t-500)+ 56.95
94.95=0.25(t-500)+ 56.95
94.95=0.25t - 125 + 56.95
94.95=0.25t - 68.05
163=0.25t
652 minutes = t
.
Upper bound:
C=0.25(t-500)+ 56.95
144.45=0.25(t-500)+ 56.95
144.45=0.25t - 125 + 56.95
144.45=0.25t - 68.05
212.5=0.25t
850 minutes = t
.
For the cost to be between $94.95 and $144.45, the telephoning time must be between 652 minutes and 850 minutes.